OpenStudy (anonymous):

Where P is (-2, 6), R is (-10, 4), and segment is a diameter of P, find the coordinates of T.

OpenStudy (anonymous):

Is that the complete question. It's not clear.

OpenStudy (anonymous):

should be segment of rt is diameter of p

OpenStudy (anonymous):

So, RT is the diameter and P is the center...P is the mid-point of RT Let T = (x,y) ( (-2+x)/2, (6+y)/2 ) = (-10,4) => (-2+x)/2 = -10 and (6+y)/2 = 4 Find x and y...

OpenStudy (anonymous):

Oh wait..

OpenStudy (anonymous):

While substituting the values, I took R as the midpoint of PT

OpenStudy (anonymous):

( (-10+x)/2, (4+y)/2 ) = (-2,6) Equate the like cordinates on both sides as I did above and find x and y..T=(x,y)